# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

### v1

@@ Abstract
Among various spatio-temporal prediction tasks, epidemic forecasting plays a critical role in public health management. Recent studies have demonstrated the strong potential of spatio-temporal graph neural networks (STGNNs) in extracting heterogeneous spatio-temporal patterns for epidemic forecasting. However, most of these methods bear an over-simplified assumption that two locations (e.g., cities) with similar observed features in previous time steps will develop similar infection numbers in the future. In fact, for any epidemic disease, there exists strong heterogeneity of its intrinsic evolution mechanisms across geolocation and time, which can eventually lead to diverged infection numbers in two “similar” locations. However, such mechanistic heterogeneity is non-trivial to be captured due to the existence of numerous influencing factors like medical resource accessibility, virus mutations, mobility patterns, etc., most of which are spatio-temporal yet unreachable or even unobservable. To address this challenge, we propose a Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN), a novel epidemic forecasting framework. By binding the epidemiology mechanistic model into a GNN, HeatGNN learns epidemiology-informed location embeddings of different locations that reflect their own transmission mechanisms over time. With the time-varying mechanistic affinity graphs computed with the epidemiology-informed location embeddings, a heterogeneous transmission graph network is designed to encode the mechanistic heterogeneity among locations, providing additional predictive signals to facilitate accurate forecasting. Experiments on three benchmark datasets have revealed that HeatGNN outperforms various strong baselines. Moreover, our efficiency analysis verifies the real-world practicality of HeatGNN on datasets of different sizes.
Index Terms:
Spatio-temporal Graphs, Predictive Analytics, Epidemic Forecasting
@@ I Introduction
The outbreak and spread of epidemics are massive disasters facing global human society, which can cause a substantial number of deaths or irreversible physiological damage to humans[1] as well as significant economic losses across various industries worldwide[2]. Recent global outbreaks, such as COVID-19, have placed unprecedented pressure on public health systems and decision-makers [3]. In response, there has been a growing interest in developing accurate and effective epidemic forecasting models[4, 5]. These models can support the implementation of public health interventions[6], optimize resource allocation[7], and enhance public awareness of early prevention and control strategies[8], thereby helping to prevent and control the onset and rapid spread of epidemics.
@@
Figure 1: Spatio-temporal and mechanistic heterogeneity between different locations in epidemic transmission. (a) Epidemic transmission with different influencing factors. (b) Example of heterogeneous epidemic transmission between two locations. (c) The analyses of two heterogeneities on a real-world dataset (US-Regions).
@@
In a nutshell, epidemic forecasting predicts the infection numbers at geographically connected regions in the future horizon based on existing records. Traditional epidemic models, like the Susceptible-Infected-Recovered (SIR) model, rely on mechanistic approaches and location networks to simulate multi-regional transmission for epidemic forecasting[9, 10, 11]. Despite advantageous explainability and applicability, these models exhibit limited capacity to capture real-world complexities. As such, with the rapid development of machine learning, methods represented by spatio-temporal graph neural networks (STGNNs) have been used to learn spatial and temporal dependencies in epidemic transmission simultaneously [5, 4, 12, 13, 14, 15, 16]. However, many STGNNs are purely data-driven methods that neglect the underlying epidemic transmission mechanisms, making them prone to overfitting historical data and being misled by spurious features. To address the limitations of both sides, epidemiological machine learning (EML) models [17, 18, 19] have been proposed by incorporating mechanistic laws into a machine learning backbone to enhance the model’s generalizability and robustness. This trend also extends to STGNNs, where the time- and location-sensitive parameters in epidemic models are predicted to support dynamic epidemic forecasting[20, 21].
@@
Generally, a key to epidemic forecasting is heterogeneity discovery, such that similar spatio-temporal patterns among locations can be aligned to enhance predictions while dissimilar ones are filtered out. This also explains the popularity of STGNNs for this task owing to their capability of learning the semantic affinity of connected locations within the topology of the graph. Meanwhile, existing models commonly use an over-simplified assumption that such heterogeneity only exists within observed features, such as past infection number, population, and demographic statistics of a region. However, in reality, epidemic heterogeneity holds for not only spatio-temporal features, but also the underlying intrinsic epidemic transmission mechanism. In this paper, we refer to these two types of heterogeneity as spatio-temporal heterogeneity and mechanistic heterogeneity, respectively. For illustration, we provide an example in Figure 1. In a nutshell, while spatio-temporal heterogeneity compares observed features (i.e., what they are now), like infection numbers, the mechanistic heterogeneity highlights the evolutionary trajectories of each location (i.e., how they will change), such as infection rates. It is worth noting that the two heterogeneity types are not mutually substitutable but rather complementary. For instance, though the two locations in Figure 1(b) have similar patterns in their infection numbers until the 20th week, the numbers have moved toward different directions in the following two weeks because of the distinct dynamics in the two locations’ infection rates. In the real epidemic transmission process, the two heterogeneity signals can vary greatly, as depicted by the Pearson Correlation Coefficient (PCC) across locations from a sampled time step (Figure 1(c)). Hence, if only the spatio-temporal heterogeneity is considered in the STGNN, the model is prone to making inaccurate predictions in the long term.
@@
To this end, bringing awareness of the mechanistic heterogeneity to the prediction model appears to be a beneficial move. However, learning such heterogeneity incurs non-trivial challenges. The first challenge lies in effectively understanding and mimicking the underlying intrinsic transmission mechanisms, which are jointly determined by a complex suite of factors. As shown in Figure 1(a), those influencing factors [22, 6, 23, 24, 25, 8, 26] include (but are not limited to) public health resources, policy interventions, climates, individual or group susceptibility, virus mutations, and so on, many of which are not linked to the epidemic records and are even unobservable. Consequently, this renders explicit and precise modeling of the intrinsic epidemic transmission mechanism infeasible. At the same time, the second challenge is the modeling of the similarity/dissimilarity across the epidemic transmission mechanism of distinct regions. The root cause is the lack of effective measures to represent and quantify the mechanistic heterogeneity across locations, preventing it from providing more predictive signals into epidemic forecasting.
@@
To address these challenges, we propose an innovative framework for heterogeneous epidemic transmission forecasting modeling: Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN). This framework aims to improve epidemic forecasting performance by simultaneously learning spatio-temporal and mechanistic heterogeneity. Firstly, we introduce epidemiology-informed embedding learning to encapsulate the mechanistic patterns into each location’s representation. To bypass the intricacy and unavailability of all possible influencing factors, instead of simulating the explicit mechanistic process of epidemic transmission, we parameterize the time-varying SIR model for each location with its spatio-temporal graph embedding. By aligning the behaviors of the implicitly parameterized time-varying SIR model and the explicitly derived one, the epidemiology-informed embedding can be effectively optimized. Secondly, to quantify and take full advantage of the mechanistic heterogeneity across locations, we propose to learn the time-varying mechanistic affinity graphs. Each mechanistic affinity graph is constructed based on the pairwise similarity between two epidemiology-informed location embeddings, where a sparsification approach is in place to remove noisy connections and ensure efficiency. With a backbone STGNN, the spatio-temporal and mechanistic heterogeneities are captured by respectively modeling the geographical and mechanistic affinity graphs across time steps to facilitate heterogeneity-aware epidemic forecasting. In summary, our main contributions are as follows:
@@
  • In the context of epidemic forecasting, we characterize two types of heterogeneity, namely, saptio-temporal heterogeneity and mechanistic heterogeneity. We point out that on top of the commonly used spatio-temporal heterogeneity, mechanistic heterogeneity is a largely overlooked yet crucial predictive signal for predicting future infection numbers.
@@
  • Our framework achieves state-of-the-art performance across three real-world epidemic datasets. Our efficiency analysis demonstrates that HeatGNN’s inference time is highly scalable w.r.t. the number of locations involved in the task.
@@
II Preliminary
II-A Problem Formalization
@@
In epidemic transmission, we formulate the epidemic forecasting problem as a node-level regression task on graphs. There are $N$ locations of interest in an area; each location (e.g., a region, city, or state) is represented as a node. Given the spatio-temporal observations of epidemic transmission risk dynamics: $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{T}\}^{T}\in\mathbb{R}^{T\times N}$, where $T$ represents the length of time step (e.g., the infected population number or infection number for $T$ weeks). Each $\boldsymbol{x}_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,N}\}\in\mathbb{R}^{N\times 1}(t=1,2,\ldots,T)$ is an $N-$dimensional vector, with $x_{t,i}(i=1,2,\ldots,N)$ being the epidemic transmission risk at the time step $t$ in the location $i$. The epidemic is transmitted in different areas, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a geographical graph, where $\mathcal{V}$ is the nodes set comprising $|\mathcal{V}|=N$ nodes (e.g., regions, cities, or states). The $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the geographic link between these nodes. The adjacency matrix $\mathcal{A}\in\mathbb{R}^{N\times N}$ of the geographical graph is formulated as $\mathcal{A}_{ij}=1$ if there is an edge $e_{ij}\in\mathcal{E}$, and $\mathcal{A}_{ij}=0$ otherwise.
@@
where $\boldsymbol{x}_{t+h}$ is an epidemic transmission risk in a future time step $t+h$ and $h$ represents the forecasting horizon (e.g., the number of time steps to be predicted ahead). $\boldsymbol{X}_{t-w+1:t}=\{\boldsymbol{x}_{t-w+1},\boldsymbol{x}_{t-w+2},\ldots,\boldsymbol{x}_{t}\}\in\mathbb{R}^{w\times N}$ is the historical epidemic transmission risk, $w$ represents the window size of historical input, and $N$ is number of locations.
III Methodologies
@@
III-A Spatio-temporal Graph Learning Module
@@
The spread of an epidemic in a city is influenced by nearby cities with similar geographic features, as adjacent locations have similar weather and topographical conditions, leading to comparable epidemic transmission patterns. Conversely, distinct weather and topography can result in varying patterns of spread. These differences lead to the emergence of spatio-temporal heterogeneity in epidemic transmission. To capture spatio-temporal heterogeneity, we propose using the STGL module in epidemic forecasting. The STGL module utilizes historical data to learn dynamic time series variations and geographic interactions, enabling it to capture the spatio-temporal heterogeneity of epidemic transmission across locations. EpiGNN is a graph neural network-based epidemic forecasting model that considers local and global spatial effects of regions and captures spatio-temporal dependencies well in some datasets[12]. Therefore, we use the backbone of EpiGNN as our STGL module. Given the historical epidemic transmission risk $\boldsymbol{X}_{t-w:t}$ and the adjacency matrix $\mathcal{A}$ of geographic topology, the spatio-temporal (ST) embedding $\boldsymbol{P}_{t}\in\mathbb{R}^{N\times D_{1}}$ with $D_{1}$-dimensional containing spatio-temporal heterogeneity information in $N$ locations can be described as:
$\boldsymbol{P}_{t}=f_{\text{ST}}(\mathcal{A},\boldsymbol{X}_{t-w:t}),$ (2)
where $f_{\text{ST}}$ is the STGL module.
III-B Epidemiology-informed Embedding Learning Module

### v2

@@ Abstract
Among various spatio-temporal prediction tasks, epidemic forecasting plays a critical role in public health management. Recent studies have demonstrated the strong potential of spatio-temporal graph neural networks (STGNNs) in extracting heterogeneous spatio-temporal patterns for epidemic forecasting. However, most of these methods bear an over-simplified assumption that two locations (e.g., cities) with similar observed features in previous time steps will develop similar infection numbers in the future. In fact, for any epidemic disease, there exists strong heterogeneity of its intrinsic evolution mechanisms across both geolocation and time, which can eventually lead to diverged infection numbers in two “similar” locations. However, capturing such mechanistic heterogeneity is non-trivial due to the existence of numerous influencing factors like medical resource accessibility, virus mutations, mobility patterns, etc., most of which are spatio-temporal yet unreachable or even unobservable. To address this challenge, we propose a Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN), a novel epidemic forecasting framework. By binding the epidemiology mechanistic model into a Graph Neural Network, HeatGNN learns epidemiology-informed location embeddings of different locations that reflect their own transmission mechanisms over time. On top of the conventional spatio-temporal module, a heterogeneous transmission graph network is designed to uncover the mechanistic heterogeneity among locations, where the time-varying mechanistic affinity graphs are computed on-the-fly and provide additional predictive signals. Experiments [^1] Code is available at https://anonymous.4open.science/r/HeatGNN-14DB. on four benchmark datasets have revealed that HeatGNN outperforms various strong baselines.
@@ I Introduction
The outbreak and spread of epidemics are massive disasters facing global human society, which can cause a substantial number of deaths or irreversible physiological damage to humans [proal2021long] as well as significant economic losses across various industries worldwide [lenzen2020global]. Recent global outbreaks, such as COVID-19, have placed unprecedented pressure on public health systems and decision-makers [adhikari2020epidemiology]. In response, there has been a growing interest in developing accurate and effective epidemic forecasting models [deng2020cola, zhan2021random]. These models can support the implementation of public health interventions [zhan2023modeling], optimize resource allocation [chen2023prediction], and enhance public awareness of early prevention and control strategies [saha2020epidemic], thereby helping to prevent and control the onset and rapid spread of epidemics.
@@
Figure 1: Spatio-temporal and mechanistic heterogeneity between different locations in epidemic transmission. (a) Epidemic transmission with different influencing factors. (b) Example of heterogeneous epidemic transmission between two locations. (c) The analyses of two heterogeneities on a dataset (US-Regions). The detailed prior experiment is provided in Section V-B.
@@
In a nutshell, epidemic forecasting predicts the infection numbers at geographically connected regions in the future horizon based on existing records. Traditional epidemic models, like the Susceptible-Infected-Recovered (SIR) model, rely on mechanistic approaches and location networks to simulate multi-regional transmission for epidemic forecasting [kermack1927contribution, zhan2021identifying]. Despite advantageous explainability and applicability, these models exhibit limited capacity to capture real-world complexities. As such, with the rapid development of machine learning, methods represented by spatio-temporal graph neural networks (STGNNs) have been used to learn spatial and temporal dependencies in epidemic transmission simultaneously [zhan2021random, deng2020cola, xie2022epignn, qiu2024msgnn, panagopoulos2021transfer, hy2022temporal, nguyen2023predicting]. However, many STGNNs are purely data-driven methods that neglect the underlying epidemic transmission mechanisms, making them prone to overfitting historical data and being misled by spurious features. To address the limitations of both sides, epidemiological machine learning models [wang2019defsi, wan2024epidemiology, rodriguez2023einns] have been proposed by incorporating mechanistic laws into a machine learning backbone to enhance the model’s generalizability and robustness. This trend also extends to STGNNs, where the time- and location-sensitive parameters in epidemic models are predicted to support dynamic epidemic forecasting [liu2023epidemiology, cao2022mepognn].
@@
Generally, a key to epidemic forecasting is heterogeneity discovery, such that similar spatio-temporal patterns among locations can be aligned to enhance predictions while dissimilar ones are filtered out. This also explains the popularity of STGNNs for this task owing to their capability of learning the semantic affinity of connected locations within the topology of the graph. Meanwhile, existing models commonly use an over-simplified assumption that such heterogeneity only exists within observed features, such as past infection number, population, and demographic statistics of a region. However, in reality, epidemic heterogeneity holds for not only spatio-temporal features but also the underlying intrinsic epidemic transmission mechanism. In this paper, we refer to these two types of heterogeneity as spatio-temporal heterogeneity and mechanistic heterogeneity, respectively. For illustration, we provide an example in Figure 1. In a nutshell, while the spatio-temporal heterogeneity refers to the differences among cities in observed features (i.e., what they are now), such as geographic location and time-dependent factors (e.g., infection numbers)., like infection numbers, the mechanistic heterogeneity highlights the evolutionary trajectories of each location (i.e., how they will change), such as infection rates (that can be described by a well-defined SIR model). It is worth noting that the two heterogeneity types are not mutually substitutable but rather complementary. For instance, though the two locations in Figure 1(b) have similar patterns in their infection numbers until the 20th week, the numbers have moved toward different directions in the following two weeks because of the distinct dynamics in the two locations’ infection rates. In the real epidemic transmission process, the two heterogeneity signals can vary greatly, as depicted by the Pearson Correlation Coefficient (PCC) across locations from a sampled time step (Figure 1(c)). Hence, if only the spatio-temporal heterogeneity is considered in the STGNN, the model is prone to making inaccurate predictions in the long term.
@@
To this end, bringing awareness of the mechanistic heterogeneity to the prediction model appears to be a beneficial move. However, learning such heterogeneity incurs non-trivial challenges. The first challenge lies in effectively understanding and mimicking the underlying intrinsic transmission mechanisms, which are jointly determined by a complex suite of factors. As shown in Figure 1(a), those influencing factors [miller2020disease, zhan2023modeling, liu2021role, davies2020age, kwok2020herd, saha2020epidemic, gibbs2020changing] include (but are not limited to) public health resources, policy interventions, climates, individual or group susceptibility, virus mutations, and so on, many of which are not linked to the epidemic records and are even unobservable. Consequently, this renders explicit and precise modeling of the intrinsic epidemic transmission mechanism infeasible. At the same time, the second challenge is the modeling of the similarity/dissimilarity across the epidemic transmission mechanisms of distinct regions. The root cause is the lack of effective measures to represent and quantify the mechanistic heterogeneity across locations, preventing it from providing more predictive signals into epidemic forecasting.
@@
To address these challenges, we propose an innovative framework for heterogeneous epidemic transmission forecasting modeling: Heterogeneous Epidemic-Aware Transmission Graph Neural Network (HeatGNN). This framework aims to improve epidemic forecasting performance by simultaneously learning spatio-temporal and mechanistic heterogeneity. Firstly, we introduce epidemiology-informed embedding learning to encapsulate the mechanistic patterns into each location’s representation. To bypass the intricacy and unavailability of all influencing factors, instead of simulating the explicit mechanistic process of epidemic transmission, we parameterize the time-varying Susceptible-Infected-Recovered (SIR) model for each location with its spatio-temporal graph embedding. By aligning the behaviors of the implicitly parameterized time-varying SIR model and the explicitly derived one, the epidemiology-informed embedding can be effectively optimized. Secondly, to quantify and take full advantage of the mechanistic heterogeneity across locations, we propose to learn the time-varying mechanistic affinity graphs. Each mechanistic affinity graph is constructed based on the pairwise similarity between two epidemiology-informed location embeddings, where a sparsification approach is in place to remove noisy connections and ensure efficiency. With a backbone STGNN, the spatio-temporal and mechanistic heterogeneities are captured by respectively modeling the geographical and mechanistic affinity graphs across time steps to facilitate heterogeneity-aware epidemic forecasting. In summary, our main contributions are as follows:
@@
  • In the context of epidemic forecasting, we characterize two types of heterogeneity, namely, spatio-temporal heterogeneity and mechanistic heterogeneity. We point out that on top of the commonly used spatio-temporal heterogeneity, mechanistic heterogeneity is a largely overlooked yet crucial predictive signal for predicting future infection numbers.
@@
  • Our framework achieves state-of-the-art performance across four real-world epidemic datasets. Our results demonstrate that HeatGNN has a high interpretability of intrinsic mechanism transmission across locations and good scalability and robustness in different scenarios.
@@
II Related Work
Epidemic Forecasting. Epidemic forecasting is a critical task in monitoring the spread of epidemics to help public health interventions quickly and avoid large-scale outbreaks [liu2024review]. The time dependence and seasonality of epidemic spread are considered at the beginning of epidemic forecasting. Time series regression models, such as ARIMA [benvenuto2020application] and VAR [shang2021regional], and machine learning models, such as random forests [zhan2021random], are widely used to solve this task, but these models cannot capture spatial dependencies and ignore the evolution of epidemic transmission. To further improve the ability of the model to capture temporal dependencies, the RNN and RNN-based models were proposed for this task [wang2020time]. CNN, Transformer, and their adopted models are used to capture multi-scale features, such as short-term and long-term temporal dependencies [wu2020deep]. Meanwhile, the emergence of Graph Neural Network (GNN) provides an effective solution to identifying spatio-temporal patterns in epidemic forecasting. In addition to most studies focusing on the impact of geographic topology [deng2020cola, panagopoulos2021transfer, xie2022epignn], some studies focus on the distinction between local and global information in epidemic spread [xie2022epignn], as well as multi-scale information [qiu2024msgnn]. However, these methods mainly focus on the spatio-temporal dependence of epidemic transmission but ignore intrinsic transmission patterns of epidemics. They also lack mechanism models to describe the transmission mechanisms behind epidemics. These limitations make it challenging to provide sufficient information to support public health decision-making.
Spatio-temporal Graph Learning. Spatio-temporal graph learning has emerged as a powerful framework for modeling complex data structures that evolve over both spatial and temporal. GCNs enable end-to-end learning of graph structures by efficiently capturing local features through parameter sharing [kipf2016semi]. GNNs later expanded from static graphs to dynamic graphs to address spatio-temporal modeling needs in dynamic network tasks [zhang2021traffic, deng2020cola]. The development of Graph Convolutional Recurrent Networks allowed dynamic graph updates in time series, capturing dynamic node relationships [seo2018structured]. To integrate temporal and spatial dependencies, STGNNs combine graph and temporal convolutions, effectively learning dynamic interactions [yu2017spatio]. In addition, advancements in STGNNs have incorporated recurrent neural networks [li2017diffusion, cini2023scalable] and Transformer-based attention [zha2024scaling] for more flexible spatio-temporal graph learning. Multiscale STGNNs [li2021multiscale] capture dependencies across different scales, while hypergraph-based STGNNs process high-order interactions [zhao2023dynamic, li2022spatial]. Graph diffusion networks model information propagation [zhang2021traffic], and heterogeneous STGNNs handle complex spatial-temporal heterogeneity [song2020spatial].
Physics-guided Machine Learning. The approximations learned by machine learning methods often ignore existing physical knowledge, leading to unexplainable conclusions [wu2024physics]. Therefore, a combination of physics and machine learning has begun to be proposed to solve the problem of lacks of interpretability in machine learning with black-box structures. A commonly used method is PINN [raissi2017physics], which is based on combining physical residual loss with data loss to distill physical epidemic knowledge into a machine learning framework for constrained learning processes while being able to effectively capture heterogeneous data in different forecasting tasks [jiang2024physics, jiang2024epidemiology, rodriguez2023einns]. Another way to combine physical models and machine learning is neural differential equations (Neural ODEs), which use the neural network as a continuous dynamic process [chen2018neural]. For example, the time-dependent states in the epidemic model are replaced by neural networks [kosma2023neural].
Some scenarios in epidemic transmission require integrating physical models into machine learning models to improve performance and interpretability. Epidemic models are based on overly simple assumptions and cannot fully understand the epidemic spread [kermack1927contribution], while machine learning models lack effective mechanisms for explanation. Network-based epidemic models capture transmission patterns between locations but face challenges in large-scale parameter optimization [zhan2021identifying] or data scarcity [gao2023evidence, cao2022mepognn]. To address those problems, some models proposed to use simulated data generated by physical models to solve the problem of insufficient data and provide high-resolution forecasting [wang2019defsi, gao2023evidence]. Meanwhile, machine learning has also begun to estimate epidemic models to solve the parameter explosion [cao2022mepognn, liu2023epidemiology]. Neural ODEs are incorporated into epidemic models to learn the continuous evolution of epidemic spread in different regions [wan2024epidemiology]. The epidemic models were used as causal mechanism models to guide the graph embedding learning process to handle noisy data and missing data better [wang2022causalgnn]. Although physics-guided machine learning has made efforts in epidemic forecasting, previous studies have ignored the mutual influence and dependency of the intrinsic epidemic transmission between locations, which makes it challenging to forecast epidemics breaking out in multiple locations.
III Problem Formalization
@@
In epidemic transmission, we formulate the epidemic forecasting problem as a node-level regression task on graphs. There are $N$ locations of interest in an area; each location (e.g., a region, city, or state) is represented as a node. Given the spatio-temporal observations of infection number: $\boldsymbol{X}=\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{T}\}^{\top}\in\mathbb{R}^{T\times N}$, where $T$ represents the length of time step (e.g., the infection number or infected population number for $T$ weeks). Each $\boldsymbol{x}_{t}=\{x_{t,1},x_{t,2},\ldots,x_{t,N}\}^{\top}\in\mathbb{R}^{N\times 1}(t=1,2,\ldots,T)$ is an NN-dimensional vector, with $x_{t,i}(i=1,2,\ldots,N)$ being the infection number at the time step $t$ in the location $i$. The epidemic is transmitted in different areas, let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a geographical graph, where $\mathcal{V}$ is the nodes set comprising $|\mathcal{V}|=N$ nodes (e.g., regions, cities, or states). The $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the geographic link between these nodes. The adjacency matrix $\mathcal{A}\in\mathbb{R}^{N\times N}$ of the geographical graph is formulated as $\mathcal{A}_{ij}=1$ if there is an edge $e_{ij}\in\mathcal{E}$, and $\mathcal{A}_{ij}=0$ otherwise.
@@
where $\boldsymbol{x}_{t+h}$ is an infection number in a future time step $t+h$ and $h$ represents the forecasting horizon (e.g., the number of time steps to be predicted ahead). $\boldsymbol{X}_{t-w+1:t}=\{\boldsymbol{x}_{t-w+1},\boldsymbol{x}_{t-w+2},\ldots,\boldsymbol{x}_{t}\}\in\mathbb{R}^{w\times N}$ is the historical infection number, $w$ represents the window size of historical input, and $N$ is the number of locations.
IV Methodologies
@@
IV-A Spatio-temporal Graph Learning Module
@@
The spread of an epidemic in a city is influenced by nearby cities with similar geographic features, as adjacent locations have similar weather and topographical conditions, leading to comparable epidemic transmission patterns. Conversely, distinct weather and topography can result in varying patterns of spread. These differences lead to the emergence of spatio-temporal heterogeneity in epidemic transmission. To capture spatio-temporal heterogeneity, we propose using the STGL module in epidemic forecasting. The STGL module utilizes historical data to learn dynamic time series variations and geographic interactions, enabling it to capture the spatio-temporal heterogeneity of epidemic transmission across locations. Given the historical infection number $\boldsymbol{X}_{t-w:t}$ and the adjacency matrix $\mathcal{A}$ of geographic topology, the spatio-temporal (ST) embedding $\boldsymbol{P}_{t}\in\mathbb{R}^{N\times D_{1}}$ with $D_{1}$-dimensional consists of the embedding $p_{t,i}$ of each location in $N$ locations can be described as:
$\boldsymbol{P}_{t}=f_{\text{ST}}(\mathcal{A},\boldsymbol{X}_{t-w+1:t}),$ (2)
where $f_{\text{ST}}$ is the STGL module. $\boldsymbol{P}_{t}$ containing spatio-temporal heterogeneity information in $N$ locations. The choice of $f_{\text{ST}}$ is general and flexible, and our HeatGNN can adapt to various spatio-temporal models. Our main innovation lies in the combination of epidemiology mechanistic models and a neural predictor. In our HeatGNN, we use the backbone of EpiGNN as our STGL module [xie2022epignn].
EpiGNN is a GNN-based model designed for epidemic forecasting, which captures spatio-temporal dependencies by integrating local and global spatial effects. It consists of three key components: multi-scale convolutions, transmission risk encoding, and the region-aware graph learner (RAGL). Multi-scale convolutions extract temporal features across different time scales, capturing both short- and long-term epidemic dynamics. For spatial transmission, EpiGNN encodes local transmission risk (LTR) via network centrality and global transmission risk (GTR) via self-attention, thereby capturing intra-regional and long-distance dependencies, respectively. Leveraging these transmission risk encodings, EpiGNN employs the RAGL to dynamically construct the region correlation graph, integrating local and global dependencies, temporal patterns, and external data. This region correlation graph informs a graph convolutional network (GCN), which aggregates information dynamically from identified neighbors. Finally, EpiGNN integrates GCN outputs with an autoregressive module to effectively capture both nonlinear spatial dependencies and linear temporal trends, thereby improving forecasting accuracy and stability.
IV-B Epidemiology-informed Embedding Learning Module